# Video Diffusion Transformers are In-Context Learners

### v1

@@ Abstract
This paper investigates a solution for enabling in-context capabilities of video diffusion transformers, with minimal tuning required for activation. Specifically, we propose a simple pipeline to leverage in-context generation: (i) concatenate videos along spacial or time dimension, (ii) jointly caption multi-scene video clips from one source, and (iii) apply task-specific fine-tuning using carefully curated small datasets. Through a series of diverse controllable tasks, we demonstrate qualitatively that existing advanced text-to-video models can effectively perform in-context generation. Notably, it allows for the creation of consistent multi-scene videos exceeding 30 seconds in duration, without additional computational overhead. Importantly, this method requires no modifications to the original models, results in high-fidelity video outputs that better align with prompt specifications and maintain role consistency. Our framework presents a valuable tool for the research community and offers critical insights for advancing product-level controllable video generation systems. The data, code, and model weights are publicly available at: https://github.com/feizc/Video-In-Context.
Machine Learning, ICML
Kunlun Inc.
@@ 1 Introduction
Diffusion models have set a new benchmark in vision generation tasks, by employing an iterative denoising process applied to a Gaussian noise latent space (Ho et al., 2020; Song et al., 2020; Peebles & Xie, 2023b; Esser et al., 2024; Rombach et al., 2022; Gu et al., 2022; Saharia et al., 2022b; Ramesh et al., 2022). Such iterative denoising process has proven to be both robust and versatile in the context of video generation. Meantime, large-scale, pre-trained video diffusion models (Blattmann et al., 2023a; Girdhar et al., 2023a; Polyak et al., 2024; Yang et al., 2024b) have greatly advanced downstream applications, with many methods providing enhanced control over various attributes, thereby enabling more precise adjustments during the generation process (Zhang & Agrawala, 2023; Ruiz et al., 2023; Gal et al., 2022; Fei et al., 2023a). Despite these advancements, adapting text-to-video models for a wide range of generative tasks, especially those requiring the creation of coherent video sets with complex intrinsic relationships, remains an open challenge (Hu & Xu, 2023; Peng et al., 2024; Zhang et al., 2023). Moreover, the generation and control of multi-scene, long-duration videos continue to pose difficulties due to the complexity of the spatio-temporal domain and the substantial GPU memory requirements.
In-context learning (ICL), initially identified as a latent capability in large autoregressive language models (Radford et al., 2021), has garnered significant attention in recent research (Dong et al., 2022). This technique involves providing models with a small set of labeled data as controllable demonstrations, enhancing their ability to reason and perform specific tasks (Brown, 2020; Dong et al., 2022; Wang et al., 2023b). ICL has been shown to notably improve model performance across a wide range of tasks (Garcia et al., 2023; Moslem et al., 2023; Qin et al., 2023; Li et al., 2023). Through ICL, models can dynamically adjust their behavior according to the contextual information provided, enabling them to execute tasks with minimal labeled data. The success of ICL in language models has sparked interest in extending these capabilities to vision tasks. However, implementing vision-based ICL presents greater challenges than in language models due to the inherently high-dimensional and diverse nature of visual data (Kirillov et al., 2019; Zhang et al., 2016), particularly when dealing with video data. As the temporal dimension in videos introduces dynamic causality and long-sequence complexity.
Inspired from (Huang et al., 2024), we aim at adapting text-to-video models for conditional control tasks with in-context capacity, offering a universal approach for multi-scene long-duration video generation. Our approach begins with a key assumption: Although trained with cut video clips, text-to-video models possess in-context generation capabilities for homologous scenes in inner connection. By effectively triggering and enhancing this built-in functionality, we aim to leverage it for more complex conditional generative tasks. Specifically, we concatenate multiple video clips from one source into an extended video sequence and combine individual video prompts into one comprehensive prompt, enabling the model to process and generate multiple-scene videos simultaneously. Then, we fine-tune a Low-Rank Adaptation (LoRA) (Hu et al., 2021) of the model with a small set of carefully designed samples. It effectively reduces the computational resources required while preserving much of the original model’s knowledge and in-context generation capabilities.
@@ 2.1 Text-to-Video Generation
Generation of video remains a longstanding challenge in the field of computer vision (Ho et al., 2022; Yu et al., 2023). In between, text-to-video generation has progressed significantly over time, evolving from early approaches based on GANs and VAEs to more advanced structures such as diffusion. Initial GAN-based methods (Vondrick et al., 2016; Saito et al., 2017; Tulyakov et al., 2018; Clark et al., 2019; Yu et al., 2022) faced dilemma with temporal coherence, leading to inconsistencies between consecutive frames. To address these issues, video diffusion models based on U-Net adapted originally designed for text-to-image, thereby improving the continuity of frames. Furthermore, DiT-based architectures (Peebles & Xie, 2023b; Lu et al., 2023; Ma et al., 2024; Gao et al., 2024) introduced spatio-temporal attention and full 3d attention, which enhanced the ability to capture the intricate dynamics of video and ensure greater frame consistency (He et al., 2022b; Blattmann et al., 2023b; Chen et al., 2023b; Girdhar et al., 2023b). Meanwhile, auto-regressive models (Yan et al., 2021; Hong et al., 2022; Villegas et al., 2022; Kondratyuk et al., 2023; Xie et al., 2024; Liu et al., 2024) employed token-based approaches to effectively model temporal dependencies, particularly excelling in the generation of long-form videos (Yin et al., 2023b; Wang et al., 2023a; Zhao et al., 2024; Henschel et al., 2024; Tan et al., 2024; Zhou et al., 2024) and in video-to-video translation tasks (Yang et al., 2023; Bao et al., 2023b; Yatim et al., 2024; Hu et al., 2024). Our work investigates the effectiveness of video diffusion Transformers for in-context capacity in long-duration multi-scene tasks and use a lightweight fine-tuning to activate it.
@@ 2.2 Diffusion Transformers
Transformer architecture (Vaswani et al., 2017) has garnered significant attention for its success in language modeling (Radford et al., 2018, 2019; Raffel et al., 2020) and substantial potential across a wide range of computer vision applications, including image classification (DOSOVITSKIY, 2020; He et al., 2022a; Touvron et al., 2021), object detection (Liu et al., 2021; Wang et al., 2021, 2022; Carion et al., 2020), and semantic segmentation (Zheng et al., 2021; Xie et al., 2021; Strudel et al., 2021), among others (Liu et al., 2022; He et al., 2022a; Fei et al., 2023c; Fei, 2021; Fei et al., 2022; Fei, 2019). Building on this success, diffusion Transformers (Peebles & Xie, 2023a) and their variants (Bao et al., 2023a; Fei et al., 2024c, e, f, d) have replaced the traditional convolutional U-Net backbone (Ronneberger et al., 2015) with Transformer-based architectures. This shift has resulted in improved scalability and simpler parameter expansion compared to U-Net-based diffusion models. Such advantages in scalability have been particularly evident in areas like video generation (Ma et al., 2024), image generation (Chen et al., 2023a), speech generation (Liu et al., 2023), and video generation (Yang et al., 2024b). Furthermore, recent innovations, such as StableAudio (Evans et al., 2024) and FluxMusic (Fei et al., 2024b, a, 2023b), have explored the application of the DiT architecture in the generation of audio and music.
@@ 2.3 Controllable Video Generation
As video generation techniques advanced, early approaches to conditioning diffusion models with spatial controls primarily relied on concatenation (Rombach et al., 2022; Saharia et al., 2022a). The emergence of large-scale, pretrained text-to-image models (Esser et al., 2021, 2024) has harnessed sophisticated text embeddings (Radford et al., 2021; Raffel et al., 2019) processed through refined attention mechanisms, propelling research into more precise control over the generation process (Li et al., 2025; Mou et al., 2024; Peng et al., 2024; Fei et al., 2023a). Notably, ControlNet (Zhang & Agrawala, 2023) facilitates fine-grained control over attributes such as depth, edges, and poses by embedding timestep-specific control features into the diffusion U-Net architecture. Additional control signals have also been explored, including camera motion (Wang et al., 2023c; He et al., 2024; Xu et al., 2024; Feng et al., 2024), objective motion (Wang et al., 2024; Yang et al., 2024a), and trajectory control (Yin et al., 2023a; Wu et al., 2025), among others (Niu et al., 2024; Fang et al., 2024). (Zhang et al., 2024) analyze the video in-context learning in autoregressive transformers. Building on (Huang et al., 2024), we propose a unified framework that consolidates these diverse controls through concatenation, generating video content using both textual prompts and masking operation.
@@ 3.2 In-Context Generation
We begin with the premise that through large-scale video pretraining, video diffusion transformer inherently understand the cross-frame video representation (Huh et al., 2024), i.e., existing text-to-video models are able to possess some level of in-context generation capability across a range of tasks, albeit with varying quality. Leveraging this insight, we aim to activate the model’s in-context generation potential through the use of meticulously curated, high-quality text-video sets. Simultaneously, we seek to maintain the integrity of the original text-to-video architecture, avoiding any structural modifications. Following (Huang et al., 2024), we propose the use of a single prompt that encompasses descriptions of multiple videos from one source, thereby enabling the desired consistent functionality within the existing framework.
Our proposed design generates a set of videos concurrently by concatenating them into a single video clip in tempera or spacial dimension during the training phase. The corresponding captions for these videos are integrated into a unified prompt template, which includes both an overarching description and specific guidance for each sub video clips. Once the video set is generated, we partition the large video into separate panels. Additionally, we apply Lora (Hu et al., 2021) to a selected subset of high-quality text-video data to enhance the capacity of models. To facilitate conditioning on an additional set of videos, we modify the inpainting scripts from image diffusion models [^1] https://huggingface.co/docs/diffusers/using-diffusers/inpaint. This training-free masking operation allows us to inpaint or outpaint the target video parts based on a partially masked video set, all of which are concatenated into the same large video.

### v2

@@ Abstract
This paper investigates a solution for enabling in-context capabilities ofmain video diffusion transformers, with minimal tuning required for activation. Specifically, we propose a simple pipeline to leverage in-context generation: (i) concatenate videos along spacial or time dimension, (ii) jointly caption multi-scene video clips from one source, and (iii) apply task-specific fine-tuning using carefully curated small datasets. Through a series of diverse controllable tasks, we demonstrate qualitatively that existing advanced text-to-video models can effectively perform in-context generation. Notably, it allows for the creation of consistent multi-scene videos exceeding 30 seconds in duration, without additional computational overhead. Importantly, this method requires no modifications to the original models, results in high-fidelity video outputs that better align with prompt specifications and maintain role consistency. Our framework presents a valuable tool for the research community and offers critical insights for advancing product-level controllable video generation systems. The data, code, and model weights are publicly available at: https://github.com/feizc/Video-In-Context.
@@ 1 Introduction
Diffusion models have set a new benchmark in vision generation tasks, by employing an iterative denoising process applied to a Gaussian noise latent space [39, 77, 61, 13, 71, 34, 75, 70]. Such iterative denoising process has proven to be both robust and versatile in the context of video generation. Meantime, large-scale, pre-trained video diffusion models [3, 32, 63, 98] have greatly advanced downstream applications, with many methods providing enhanced control over various attributes, thereby enabling more precise adjustments during the generation process [104, 73, 29, 19]. Despite these advancements, adapting text-to-video models for a wide range of generative tasks, especially those requiring the creation of coherent video sets with complex intrinsic relationships, remains an open challenge [44, 62, 107]. Moreover, the generation and control of multi-scene, long-duration videos continue to pose difficulties due to the complexity of the spatio-temporal domain and the substantial GPU memory requirements.
@@
In-context learning (ICL), initially identified as a latent capability in large autoregressive language models [67], has garnered significant attention in recent research [10]. This technique involves providing models with a small set of labeled data as controllable demonstrations, enhancing their ability to reason and perform specific tasks [5, 10, 87]. ICL has been shown to notably improve model performance across a wide range of tasks [31, 57, 64, 50]. Through ICL, models can dynamically adjust their behavior according to the contextual information provided, enabling them to execute tasks with minimal labeled data. The success of ICL in language models has sparked interest in extending these capabilities to vision tasks. However, implementing vision-based ICL presents greater challenges than in language models due to the inherently high-dimensional and diverse nature of visual data [47, 105], particularly when dealing with video data. As the temporal dimension in videos introduces dynamic causality and long-sequence complexity.
Inspired from [45], we aim at adapting text-to-video models for conditional control tasks with in-context capacity, offering a universal approach for multi-scene long-duration video generation. Our approach begins with a key assumption: Although trained with cut video clips, text-to-video models possess in-context generation capabilities for homologous scenes in inner connection. By effectively triggering and enhancing this built-in functionality, we aim to leverage it for more complex conditional generative tasks. Specifically, we concatenate multiple video clips from one source into an extended video sequence and combine individual video prompts into one comprehensive prompt, enabling the model to process and generate multiple-scene videos simultaneously. Then, we fine-tune a Low-Rank Adaptation (LoRA) [42] of the model with a small set of carefully designed samples. It effectively reduces the computational resources required while preserving much of the original model’s knowledge and in-context generation capabilities.
@@ 2.1 Text-to-Video Generation
Generation of video remains a longstanding challenge in the field of computer vision [40, 103]. In between, text-to-video generation has progressed significantly over time, evolving from early approaches based on GANs and VAEs to more advanced structures such as diffusion. Initial GAN-based methods [84, 76, 81, 9, 102] faced dilemma with temporal coherence, leading to inconsistencies between consecutive frames. To address these issues, video diffusion models based on U-Net adapted originally designed for text-to-image, thereby improving the continuity of frames. Furthermore, DiT-based architectures [61, 55, 56, 30] introduced spatio-temporal attention and full 3d attention, which enhanced the ability to capture the intricate dynamics of video and ensure greater frame consistency [37, 4, 8, 33]. Meanwhile, auto-regressive models [95, 41, 83, 48, 93, 51] employed token-based approaches to effectively model temporal dependencies, particularly excelling in the generation of long-form videos [101, 85, 108, 38, 79, 110] and in video-to-video translation tasks [97, 2, 99, 43]. Our work investigates the effectiveness of video diffusion Transformers for in-context capacity in long-duration multi-scene tasks and use a lightweight fine-tuning to activate it.
@@ 2.2 Diffusion Transformers
Transformer architecture [82] has garnered significant attention for its success in language modeling [65, 66, 69] and substantial potential across a wide range of computer vision applications, including image classification [11, 36, 80], object detection [53, 88, 89, 6], and semantic segmentation [109, 92, 78], among others [54, 36, 21, 17, 18, 16]. Building on this success, diffusion Transformers [60] and their variants [1, 24, 26, 27, 25] have replaced the traditional convolutional U-Net backbone [72] with Transformer-based architectures. This shift has resulted in improved scalability and simpler parameter expansion compared to U-Net-based diffusion models. Such advantages in scalability have been particularly evident in areas like video generation [56], image generation [7], speech generation [52], and video generation [98]. Furthermore, recent innovations, such as StableAudio [14] and FluxMusic [23, 22, 20], have explored the application of the DiT architecture in the generation of audio and music.
@@ 2.3 Controllable Video Generation
As video generation techniques advanced, early approaches to conditioning diffusion models with spatial controls primarily relied on concatenation [71, 74]. The emergence of large-scale, pretrained text-to-image models [12, 13] has harnessed sophisticated text embeddings [67, 68] processed through refined attention mechanisms, propelling research into more precise control over the generation process [49, 58, 62, 19]. Notably, ControlNet [104] facilitates fine-grained control over attributes such as depth, edges, and poses by embedding timestep-specific control features into the diffusion U-Net architecture. Additional control signals have also been explored, including camera motion [90, 35, 94, 28], objective motion [86, 96], and trajectory control [100, 91], among others [59, 15]. [106] analyze the video in-context learning in autoregressive transformers. Building on [45], we propose a unified framework that consolidates these diverse controls through concatenation, generating video content using both textual prompts and masking operation.
@@ 3.2 In-Context Generation
We begin with the premise that through large-scale video pretraining, video diffusion transformer inherently understand the cross-frame video representation [46], i.e., existing text-to-video models are able to possess some level of in-context generation capability across a range of tasks, albeit with varying quality. Leveraging this insight, we aim to activate the model’s in-context generation potential through the use of meticulously curated, high-quality text-video sets. Simultaneously, we seek to maintain the integrity of the original text-to-video architecture, avoiding any structural modifications. Following [45], we propose the use of a single prompt that encompasses descriptions of multiple videos from one source, thereby enabling the desired consistent functionality within the existing framework.
Our proposed design generates a set of videos concurrently by concatenating them into a single video clip in tempera or spacial dimension during the training phase. The corresponding captions for these videos are integrated into a unified prompt template, which includes both an overarching description and specific guidance for each sub video clips. Once the video set is generated, we partition the large video into separate panels. Additionally, we apply Lora [42] to a selected subset of high-quality text-video data to enhance the capacity of models. To facilitate conditioning on an additional set of videos, we modify the inpainting scripts from image diffusion models [^1] https://huggingface.co/docs/diffusers/using-diffusers/inpaint. This training-free masking operation allows us to inpaint or outpaint the target video parts based on a partially masked video set, all of which are concatenated into the same large video.